Author’s Guide for ICARCV2002

C. C. Cheah

School of Electrical and Electronic Engineering Nanyang Technological University Block S1, Nanyang Avenue, Singapore 639798 Email: icarcv2002@ntu.edu.sg

Abstract

The proceedings of the Seventh International Conference on Control, Automation, Robotics and Vision will be published on a CD-ROM. Manuscripts are to be submitted in electronic format as Postscript (PS) or Adobe Portable Document Format (PDF) files, preferably with a resolution no less than 600 dpi. To minimise the risk of printing errors, for PDF files, only type 1 fonts should be used. No page number should appear in the final manuscript.

1 Introduction

In order to ensure that a high quality conference CD-ROM will be produced, authors are advised to prepare their manuscripts following the guidelines detailed in this guide. Six pages are allowed for each paper. Up to two additional pages will be permitted for a charge of $ 100 for each additional page. Figures and illustrations are included in the page count. The layout of this guide can be served as a sample. Should there be any doubt, please contact us via email: ICARCV2002@ntu.edu.sg.

2 General Guidelines

Authors should prepare their manuscripts in double column format for printing on A4 (210mm x 297mm) paper, using Times Roman 10pt typeface, and must ensure that the typing area (centred) matches the guidelines given below:

(a) No page number should appear in the final manuscript.

(b) On the first page, the distance from the top edge of the paper to the top of the title should be 30mm.

(c) On the second and subsequent pages, the distance from the top edge of the paper to the top line of type should be 25mm.

(d) The bottom margin should be 25mm.

(e) The left and right margins must be the same (20mm).

(f) The width of each column should be 80mm and the separation between the two columns should be 10mm.

3 More Detailed Guidelines

Section headings should be in capital letters, bold font and numbered consecutively, starting with Introduction. Sub-section headings should be numbered 1.1, 1.2 etc. Footnotes should be avoided if possible.

3.1 Guidelines for Figures

The figure number and caption should be typed below the illustration, left justified, with subsequent lines indented.

3.2 Guidelines for References

List and number all references at the end of the paper. When citing references in the text, type the corresponding number in square brackets as shown at the end of this sentence [1].

4 More Detailed Guidelines

Section headings should be in capital letters, bold font and numbered consecutively, starting with Introduction. Sub-section headings should be numbered 1.1, 1.2 etc. Footnotes should be avoided if possible.

4.1 Guidelines for Figures

The figure number and caption should be typed below the illustration, left justified, with subsequent lines indented.

4.2 Guidelines for References

List and number all references at the end of the paper. When citing references in the text, type the corresponding number in square brackets as shown at the end of this sentence [1].

5 More Detailed Guidelines

Section headings should be in capital letters, bold font and numbered consecutively, starting with Introduction. Sub-section headings should be numbered 1.1, 1.2 etc. Footnotes should be avoided if possible.

5.1 Guidelines for Figures

The figure number and caption should be typed below the illustration, left justified, with subsequent lines indented.

5.2 Guidelines for References

List and number all references at the end of the paper. When citing references in the text, type the corresponding number in square brackets as shown at the end of this sentence [1].

6 More Detailed Guidelines

Section headings should be in capital letters, bold font and numbered consecutively, starting with Introduction. Sub-section headings should be numbered 1.1, 1.2 etc. Footnotes should be avoided if possible.

6.1 Guidelines for Figures

The figure number and caption should be typed below the illustration, left justified, with subsequent lines indented.

6.2 Guidelines for References

List and number all references at the end of the paper. When citing references in the text, type the corresponding number in square brackets as shown at the end of this sentence [1].

7 More Detailed Guidelines

Section headings should be in capital letters, bold font and numbered consecutively, starting with Introduction. Sub-section headings should be numbered 1.1, 1.2 etc. Footnotes should be avoided if possible.

7.1 Guidelines for Figures

The figure number and caption should be typed below the illustration, left justified, with subsequent lines indented.

7.2 Guidelines for References

List and number all references at the end of the paper. When citing references in the text, type the corresponding number in square brackets as shown at the end of this sentence [1].

8 Conclusions

Thank you for your co-operation and contribution. We look forward to meeting you at the ICARCV2002 in Singapore.

References

[1] X. Author, “Title of Paper,” Int. Journal of Robotic System, vol. 12, no. 2, pp. 119–134, 1999.

[2] Y. Author, Title of Book. New York, ABS Press, 1988, Chap. 6, pp.23-35.

[3] C. C. Cheah, S. Kawamura, and S. Arimoto, “Title of Paper,” in Proceedings of IEEE Int. Conference on Robotics and Automation, (Leuven, Belgium), pp. 3607–3612, 1998.